OpenStudy (anonymous):

-4/x+3/x^2=-1 help me find x please

OpenStudy (espex):

What happens if you multiply each side by x^2?

OpenStudy (espex):

\[x^{2}(\frac{-4}{x}) +(x^{2})(\frac{3}{x^{2}})=x^{2}(-1) \]